He CLUSTERS AS A MATRIX FOR PREPARATION OF Na-(H_{2})_{n} COMPLEXES

\underline{C. CALLEGARI}, J. HIGGINS, J. REHO, K. K. LEHMANN, and G. SCOLES, Chemistry Department, Princeton University, Princeton, NJ 08544.

Large He clusters (containing up to 10^{4} atoms) have been employed as a substrate to produce Na-(H_{2})_{n} (n < 10) complexes, making use of the pick-up technique. Laser Induced Fluorescence and Beam Depletion are used to spectroscopically investigate such complexes in the frequency region of Na D lines. A strong exponential decay of the total amount of collected fluorescence is observed as a function of H_{2} pressure in the pick-up region, while Beam Depletion Spectra intensities are virtually unaffected, thus demonstrating the high quenching efficiency of H_{2}. Spectra are then compared to those of Na atoms on pure He or H_{2} clusters. Interestingly, spectra of Na-(H_{2})_{n} complexes show spectral features similar to those of Na on pure H_{2} clusters only with smaller shifts with respect to the D lines position. Time resolved measurements are currently being taken and will be presented at the meeting.